Math 331 Transformation Geometry Assignments

Spring Semester, 2004

PDF and DVI (requires TeX^{1} software) versions of this page are available for printing. Most of these assignments are simply exercises designed to prepare you for the weekly quizzes. You may find it helpful to discuss these exercises with others, but if you watch someone else do an exercise, you will likely lose its benefit unless you subsequently work through the exercise privately. It is generally not helpful to seek explanation of an exercise that you have not attempted. When you have trouble with an exercise, it is important to try to state in words exactly where in that exercise and in what way you find yourself blocked. Assignments are listed by the date due.

Wed May 12

Final Examination 3:30 – 5:30 p.m. Tue May 11

Office hours: 2:30 – 3:30 p.m. Mon May 10

Office hours: 2:30 – 3:30 p.m. Fri May 7

Office hours: 2:30 – 3:30 p.m. Wed May 5

Review Session: bring questions

Expect a Quiz Mon May 3

Review Session: bring questions Fri Apr 30 (tg040430.html, also available as PDF or DVI) Wed Apr 28 (tg040428.html, also available as PDF or DVI) Mon Apr 26 (tg040426.html, also available as PDF or DVI) Fri Apr 23 (tg040423.html, also available as PDF or DVI) Wed Apr 21 (tg040421.html, also available as PDF or DVI) Mon Apr 19 (tg040419.html, also available as PDF or DVI) Fri Apr 16 (tg040416.html, also available as PDF or DVI)

1URI: http://www.tug.org/

Wed Apr 14 (tg040414.html, also available as PDF or DVI) Mon Apr 12 (tg040412.html, also available as PDF or DVI) Fri Apr 2 (tg040402.html, also available as PDF or DVI) Wed Mar 31 (tg040331.html, also available as PDF or DVI) Mon Mar 29 (tg040329.html, also available as PDF or DVI) Fri Mar 26

No assignment; but, depending on class performance on the midterm, there may be a quiz either Friday or Monday the 29th. It is suggested that you revisit questions on the midterm, if any, that you did not know.

Wed Mar 24

Midterm test (during the regular class period) Mon Mar 22

Review Session: Bring questions Fri Mar 19 (tg040319.html, also available as PDF or DVI)

The quiz announced for Wednesday will be on Friday.

See the comments on exercises^{2} (also available as PDF^{3} or DVI^{4}) regarding lines stabilized by affine transformations.

Wed Mar 17 (tg040317.html, also available as PDF or DVI) Mon Mar 15 (tg040315.html, also available as PDF or DVI) Fri Mar 12 (tg040312.html, also available as PDF or DVI) Wed Mar 10 (tg040310.html, also available as PDF or DVI) Mon Mar 8 (tg040308.html, also available as PDF or DVI) Fri Mar 5 (tg040305.html, also available as PDF or DVI) Wed Mar 3 (tg040303.html, also available as PDF or DVI)

Expect a quiz. Mon Mar 1 (tg040301.html, also available as PDF or DVI) Fri Feb 27 (tg040227.html, also available as PDF or DVI) Wed Feb 25 (tg040225.html, also available as PDF or DVI) Mon Feb 23 (tg040223.html, also available as PDF or DVI)

2URI: tg040308so.html 3URI: tg040308so.pdf 4URI: tg040308so.dvi

Fri Feb 20 (tg040220.html, also available as PDF or DVI) Wed Feb 18 (tg040218.html, also available as PDF or DVI) Fri Feb 13 (tg040213.html, also available as PDF or DVI) Wed Feb 11 (tg040211.html, also available as PDF or DVI) Mon Feb 9

Use the theorem on the sheet for the Feb. 2 assignment^{5} to prove that the three altitudes of a triangle intersect in a point. From those considerations find a triple (u, v, w) of numbers depending on the sides and the angles in the triangle that is proportional to the triple of barycentric coordinates of the altitude intersection point relative to the three vertices of the triangle.

Fri Feb 6 No Class. Friday classes were suspended by the University from 12:20 due to icy traffic conditions

Wed Feb 4 (tg040204.html, also available as PDF or DVI) Mon Feb 2 (tg040202.html, also available as PDF or DVI) Fri Jan 30 (tg040130.html, also available as PDF or DVI) Wed Jan 28 (tg040128.html, also available as PDF or DVI) Mon Jan 26 (tg040126.html, also available as PDF or DVI) Fri Jan 23 (tg040123.html, also available as PDF or DVI)

UP | TOP | Department Document network location for HTML:

http://math.albany.edu:8000/math/pers/hammond/course/mat331s2004/assgt/gassgt331.html

5URI: tg040202.html#generalceva